Problem 3 Suppose that Y. whose density is that Y Y sdom smumph from Y, is a random sample from a population f(10) ii. 1-0(y-1/2),0 < y< 1 where 0 0 <1 is a parameter. I te Caleulate the expectation of T
Answer #1

E[Y^2] = \int_{0}^{1} y^2(1 - \theta (y - 1/2))~dy = \int_{0}^{1} y^2 - \theta (y^3 - y^2/2)~dy

= \left [ y^3/3 - \theta (y^4/4 - y^3/6) \right ]_{0}^{1} = 1/3 - \theta (1/4 - 1/6) = 1/3 - \theta /12

Expectation of T is,

E[T] = E\left [ 1 - (3/n) \sum_{i=1}^{n}Y_i^2 \right ] = 1 - 3 E\left [\sum_{i=1}^{n}Y_i^2/n \right ] = 1 - 3E[Y^2]

= 1 - 3\left ( 1/3 - \theta /12 \right ) = 1 - 1 + \theta /4 = \theta /4